OpenStudy (anonymous):

Your piggy bank has a total of 47 coins in it; some are dimes and some are nickels. If you have a total of $3.95, how many nickels do you have?

OpenStudy (anonymous):

Help Please.. :)

OpenStudy (amistre64):

set up 2 equations in 2 unknowns the number of dimes and nickels = 47: a + b = 47 some set of "a" dimes and "b" nickels = 3.95: a(.10) + b(.05) = 3.95

OpenStudy (amistre64):

systems of equations can be solved by substitution, or elimination

OpenStudy (amistre64):

a + b = 47 a(.10) + b(.05) = 3.95 pick your poison :)

OpenStudy (amistre64):

since we want to get the number of nickels (b) id opt to get rid of the "a" parts

OpenStudy (anonymous):

im confused lol

OpenStudy (anonymous):

Given: d = qty of dimes Given: n = qty of nickels Given: d + n = total coins =47 Given: total $ = 3.95 Write equations: d + n = 47 10d + 5n = $3.95 d = 47 - n substitute: 10(47 - n) + 5n = 395 470 - 10n + 5n = 395 5n = 75 n = 15 ea d = 47 - 15 = 32 ea CHECK: (10)(32) + (5)(15) = 395/100 = $3.95

OpenStudy (anonymous):

is this better ?

OpenStudy (anonymous):

yes thank you :)

OpenStudy (anonymous):

yur welcome =)